# AutoHFormer: Efficient Hierarchical Autoregressive Transformer for Time Series Prediction

### v1

@@ Abstract
Time series forecasting requires architectures that simultaneously achieve three competing objectives: (1) strict temporal causality for reliable predictions, (2) sub-quadratic complexity for practical scalability, and (3) multi-scale pattern recognition for accurate long-horizon forecasting. We introduce AutoHFormer, a hierarchical autoregressive transformer that addresses these challenges through three key innovations: 1) Hierarchical Temporal Modeling: Our architecture decomposes predictions into segment-level blocks processed in parallel, followed by intra-segment sequential refinement. This dual-scale approach maintains temporal coherence while enabling efficient computation. 2) Dynamic Windowed Attention: The attention mechanism employs learnable causal windows with exponential decay, reducing complexity while preserving precise temporal relationships. This design avoids both the anti-causal violations of standard transformers and the sequential bottlenecks of RNN hybrids. 3) Adaptive Temporal Encoding: a novel position encoding system is adopted to capture time patterns at multiple scales. It combines fixed oscillating patterns for short-term variations with learnable decay rates for long-term trends. Comprehensive experiments demonstrate that AutoHFormer 10.76× faster training and 6.06× memory reduction compared to PatchTST on PEMS08, while maintaining consistent accuracy across 96-720 step horizons in most of cases. These breakthroughs establish new benchmarks for efficient and precise time series modeling. Implementations of our method and all baselines in hierarchical autoregressive mechanism are available at https://github.com/lizzyhku/Autotime.
@@ I Introduction
Time series forecasting [1, 2, 3, 4, 5] stands as a fundamental pillar of modern predictive analytics, enabling data-driven decision making across numerous mission-critical domains. As demonstrated in recent literature [6], this task has become increasingly vital in our data-rich era. In financial markets [7, 8], accurate forecasting drives risk management strategies. Energy sector applications require precise demand predictions to optimize renewable integration. Transportation systems [9, 10] rely on temporal pattern analysis for traffic flow optimization. The importance of time series analysis also extends to public safety through crime prediction [6]. These diverse applications share a common need: the ability to extract meaningful patterns from temporal data to inform proactive decision-making.
@@
The evolution of time series forecasting has been significantly advanced by Transformer architectures [11], yet persistent limitations hinder their practical deployment. As shown in Figure 1, conventional Transformers suffer from anti-causal attention flows, violating the fundamental $p(x_{t}|x_{<t})$ principle through bidirectional information propagation, which is particularly detrimental in autoregressive settings. Simultaneously, RNN-Transformer hybrids impose sequential computation bottlenecks ($\mathcal{O}(L)$ time steps) ($L$ is the sequence length) that eliminate parallel processing advantages. Three fundamental shortcomings emerge from this architectural landscape: (1) Quadratic Complexity: Standard self-attention’s $\mathcal{O}(L^{2})$ memory and computation requirements render long-sequence processing infeasible; (2) Error Propagation: Autoregressive approaches accumulate prediction errors exponentially across time steps; (3) Temporal Representation Rigidity: Fixed attention patterns fail to adapt to varying time scales and event frequencies.
While recent innovations like Informer [12] and Pyraformer [13] address individual aspects, they fail to address three competing demands simultaneously: computational efficiency (sub-quadratic complexity), temporal coherence (strict causality), and multi-scale pattern capture (from minute-level transients to daily periodicity). To resolve this trilemma, we propose AutoHFormer, a novel architecture that addresses three critical limitations of existing approaches via three innovations: (1) Sub-quadratic Complexity and Temporal Coherence Maintenance. The Dynamic Windowed Attention Mechanism achieves $\mathcal{O}(LW)$ ($W$ denotes the window size) complexity via learnable causal windows, maintaining temporal coherence through: (i) strict architectural causality ($p(x_{t}|x_{<t})$), (ii) exponential decay ($\tau(t,t^{\prime})=e^{-\gamma|t-t^{\prime}|}$) to weight dependencies by proximity (shown in Figure 1). This design ensures sub-quadratic efficiency ($\lim_{L\to\infty}LW/L^{2}=0$) while preserving parallelizability. The constant memory footprint ($\mathcal{O}(W)$ per token) enables efficient deployment without sacrificing temporal relationships. (2) Error-Resistant Hierarchical Refinement. Autoregressive approaches suffer from compounding prediction errors. Motivated by this, we propose dual-phase architecture with (i) segment-level autoregression and (ii) intra-segment stepwise correction via normalized residuals. Meanwhile, for inter-segment level autoregression, layer-normalized residual updating and moving average smoothing are proposed to enable stable prediction within segment. It not only reduces error propagation specifically in long-term forecasts, but also provides stable long-term prediction. (3) Adaptive Temporal Position Encoding. Fixed positional representations fail to capture varying temporal scales. Hybrid system combining sinusoidal patterns with learned decay coefficients. It captures both local dynamics through high-frequency sinusoidal components and global trends via learned decay rates. The synergistic integration of these components establishes AutoHFormer as a new paradigm in temporal modeling, achieving better prediction performance in comparable complexity.
@@
  • An Efficient and Unified Hierarchical Autoregressive Transformer for Time Series Prediction. We propose a hierarchal autoregressive framework via block (segment) processing and step refinement for efficient long-horizon prediction.
  • Efficient Dynamic Attention with Temporal Coherence Maintenance. A memory-efficient attention mechanism featuring is implemented with Learnable causal windows and Exponential decay.
  • Adaptive Temporal Encoding for Multi-scale Patterns and Short Transients. We adopts relative positive temporal encoding with learned decay coefficients, which captures varying temporal scales including short transients and multi-scale patterns.
  • Comprehensive Empirical Validation. Through extensive experiments on benchmark datasets, AutoHFormer demonstrates superior performance compared to 8 state-of-the-art baselines, including transformer variants (iTransformer, Informer, Autoformer) and modern architectures (PatchTST, TimeMixer), particularly excelling in long-sequence scenarios (720+ prediction steps). Implementations of our method and all baselines in hierarchical autoregressive mechanism are available at https://github.com/lizzyhku/Autotime.
@@ II-A Overview of AutoHFormer
The AutoHFormer framework introduces a novel hierarchical autoregressive transformer that fundamentally advances time series forecasting through three key innovations. First, our hierarchical architecture combines segment-wise processing with stepwise refinement to eliminate error propagation in long-horizon predictions. Second, the optimized windowed attention mechanism employs learnable causal windows $\mathcal{W}_{t}=[t-w,t]$ ($W$: window size) to achieve $\mathcal{O}(LW)$ complexity while preserving temporal coherence ($p(x_{t}|x_{<t})$). The resulting system maintains sub-quadratic scaling ($LW\ll L^{2}$ for $L\gg W$) with full parallelizability, while the fixed per-token memory requirement ($\mathcal{O}(W)$) ensures hardware efficiency. Third, adaptive temporal position encodings capture multi-scale patterns and short transients through hybrid sinusoidal-decay embeddings. Theoretically grounded in Section II-H and empirically validated in Section III, AutoHFormer demonstrates: (1) 10.76$\times$ faster training than conventional transformer-based PatchTST on PEMS08, (2) consistent accuracy across 96-720 step horizons in most of cases, and (3) 6.06$\times$ memory reduction than PatchTST on PEMS08 establishing new standards for efficient and accurate temporal modeling.
@@ II-B Preliminary
For a time series with $T$ observations, we can express the autoregressive mechanism in matrix form:
$\begin{bmatrix}X_{p+1}\\ X_{p+2}\\ \vdots\\ X_{T}\end{bmatrix}=\Omega\begin{bmatrix}c\\ \phi_{1}\\ \phi_{2}\\ \vdots\\ \phi_{p}\end{bmatrix}+\begin{bmatrix}\epsilon_{p+1}\\ \epsilon_{p+2}\\ \vdots\\ \epsilon_{T}\end{bmatrix}$ (2)
where $\Omega=\begin{bmatrix}1&X_{p}&X_{p-1}&\cdots&X_{1}\\ 1&X_{p+1}&X_{p}&\cdots&X_{2}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&X_{T-1}&X_{T-2}&\cdots&X_{T-p}\end{bmatrix}$. Here Equation 2 contains lagged values of $X_{t}$ and enables efficient computation of parameters via least squares. Menahwile, $T>2p$ is required for identifiability. An AR process is stationary if all roots of the characteristic polynomial lie outside the unit circle:
$1-\phi_{1}z^{-1}-\phi_{2}z^{-2}-\cdots-\phi_{p}z^{-p}\neq 0\quad\text{for}\quad|z|\leq 1$ (3)
@@ II-B2 Problem Definition
$\widehat{y}_{(h-1)H+t}=\mathcal{G}\big{(}\mathbf{X}_{1:L},\widehat{\mathbf{Y}}_{1:h-1},\widehat{y}_{(h-1)H+1},...,$ (5)
$\widehat{y}_{(h-1)H+t-1}\big{)}\quad\forall t\in\{1,...,H\}$
@@ II-C Hierarchical Autoregressive Generative Mechanism
    $\widehat{y}_{(h-1)H+t}=\mathcal{G}_{\phi}(\mathbf{C}_{h},\widehat{y}_{(h-1)H+1},...,\widehat{y}_{(h-1)H+t-1})$ (9)
    $\quad\forall t\in\{1,...,H\}$
@@
[…]
@@ II-D1 Dynamic Windowed Masked Attention (DWMA)
Time series forecasting demands efficient modeling of hierarchical temporal patterns, where both fine-grained fluctuations and coarse-grained trends jointly determine future states. Existing attention mechanisms [11] fundamentally struggle with this multi-scale requirement due to either quadratic complexity ($\mathcal{O}(L^{2})$) or information loss in simplified approximations. Our DWMA breakthrough addresses these limitations via: i) Adaptively Constrained Attention Windows: The designation of adaptively constrained attention window is shown as $\mathcal{W}_{t}=\left[\max\left(1,t-\frac{W}{2}\right),t\right]$, which establishes dynamic receptive fields that maintain strict causality while enabling $\mathcal{O}(LW)$ computational complexity. The window size $W$ acts as a hyperparameter controlling the trade-off between context range and efficiency. ii) Learnable Exponential Decay: It introduces position-sensitive weighting via $\tau(t,t^{\prime})=\exp\left(-\gamma|t-t^{\prime}|\right)$, where the trainable parameter $\gamma$ automatically adapts to dataset characteristics. This creates continuous attention spectra that emphasize recent patterns while preserving access to critical long-range dependencies.
DWMA’s hybrid design achieves superior efficiency-accuracy trade-offs compared to vanilla attention [11]. The decay mechanism automatically adjusts to diverse temporal patterns, which enables our model to have better performance in time series datasets.
@@ II-D3 Obtaining Embeddings Based on Enhanced Attention
In former steps, the attention weights are computed by incorporating the Precomputed Position Encodings (PPE). For each time step $t$ and its neighboring time steps $t^{\prime}$ within the sliding window of size $W$, the attention weights $A_{t,t^{\prime}}$ are calculated as above Equation 13. The attention weights $A_{t,t^{\prime}}$ are then used to compute the hidden states $\mathbf{H}_{\text{transformer}}$ for each time step $t$:
$\mathbf{H}_{\text{transformer},t}=\sum_{t^{\prime}=1}^{L}A_{t,t^{\prime}}\cdot V_{t^{\prime}},$ (14)
Here $V_{t^{\prime}}\in\mathbb{R}^{d_{v}}$ is the value vector for time step $t^{\prime}$ and $d_{v}$ is the value dimension. The resulting hidden states $\mathbf{H}_{\text{transformer}}\in\mathbb{R}^{L\times d}$ capture the temporal dependencies in the input time series data, enriched by the relative position encodings.
@@ II-E Model Optimization
The training objective combines prediction accuracy with architectural constraints tailored for autoregressive time series modeling (Algorithm 1) via Equation 6.
Adam updates with gradient clipping (max norm 1.0) and learning rate $\eta=10^{-4}$. The moving average update minimizes error accumulation through $\widehat{y}_{(h-1)H+t}\leftarrow W_{o}\mathbf{f}_{t}[-1]$.
@@ II-F Model Complexity
Our AutoHFormer architecture achieves efficient resource utilization through two key mechanisms (Algorithm 1, Lines 10-13): First, the windowed attention reduces time complexity from $\mathcal{O}(L^{2}d)$ to $\mathcal{O}(LWd)$ by processing only local contexts $\mathcal{W}_{t}=[\max(1,t-W/2),t]$. As shown in Table 2, this yields 32$\times$ faster computation than full attention when $L=1024$ and $W=32$. Second, the autoregressive loop (Lines 6-20) maintains $\mathcal{O}(Wd)$ memory overhead through incremental prediction cache updates, avoiding the $\mathcal{O}(L^{2})$ memory bottleneck of conventional transformers. The dominant terms simplify to:
@@
Comparative Advantages: AutoHFormer is able to maintain sub-quadratic scaling in both time and memory versus quadratic growth in vanilla attention. Three architectural innovations enable this: (1) The dynamic masked attention mechanism enables it to deal with adaptive sequence lengths, (2) Relative position embeddings $p_{t-t^{\prime}}$ require only $\mathcal{O}(Wd)$ storage, and (3) LayerNorm residuals allow stable depth scaling. The result is 64$\times$ higher than RNNs at $L=2048$ while preserving transformer-grade accuracy (Table III).
@@ II-G Algorithm of AutoHFormer
5:$\mathbf{X}_{\text{proj}}\leftarrow\text{Linear}(d)(\mathbf{X})$triangleright Project input to $d$-dim space
@@
8:     $\mathbf{C}_{h}\leftarrow\text{Concat}(\mathbf{X},\widehat{\mathbf{Y}})$triangleright Global context
9:     $\mathbf{M}_{h}\leftarrow\text{SegmentMask}(\mathbf{C}_{h},H)$triangleright Segment causality mask
@@
11:     $\widehat{\mathbf{Y}}_{h}^{init}\leftarrow\mathcal{F}_{\theta}(\mathbf{C}_{h})\in\mathbb{R}^{H\times d}$triangleright Initial segment prediction
@@
15:         $\mathcal{W}_{t}\leftarrow\{t^{\prime}|\max(1,t-W/2)\leq t^{\prime}\leq t\}$triangleright Local causal window
@@
17:         $\tau_{t}\leftarrow\exp(-\gamma\cdot|t-\mathcal{W}_{t}|)$triangleright Adaptive decay
@@
22:         $\widehat{y}_{(h-1)H+t}\leftarrow W_{o}\mathbf{f}_{t}[-1]$triangleright Final prediction
@@
24:     $\widehat{\mathbf{Y}}\leftarrow\text{Concat}(\widehat{\mathbf{Y}},\widehat{\mathbf{Y}}_{h})$triangleright Aggregate segments

### v2

@@ Abstract
Time series forecasting requires architectures that simultaneously achieve three competing objectives: (1) strict temporal causality for reliable predictions, (2) sub-quadratic complexity for practical scalability, and (3) multi-scale pattern recognition for accurate long-horizon forecasting. We introduce AutoHFormer, a hierarchical autoregressive transformer that addresses these challenges through three key innovations: 1) Hierarchical Temporal Modeling: Our architecture decomposes predictions into segment-level blocks processed in parallel, followed by intra-segment sequential refinement. This dual-scale approach maintains temporal coherence while enabling efficient computation. 2) Dynamic Windowed Attention: The attention mechanism employs learnable causal windows with exponential decay, reducing complexity while preserving precise temporal relationships. This design avoids both the anti-causal violations of standard transformers and the sequential bottlenecks of RNN hybrids. 3) Adaptive Temporal Encoding: a novel position encoding system is adopted to capture time patterns at multiple scales. It combines fixed oscillating patterns for short-term variations with learnable decay rates for long-term trends. Comprehensive experiments demonstrate that AutoHFormer 10.76× faster training and 6.06× memory reduction compared to PatchTST on PEMS08, while maintaining consistent accuracy across 96-720 step horizons in most of cases. These breakthroughs establish new benchmarks for efficient and precise time series modeling. Implementations of our method and all baselines in hierarchical autoregressive mechanism are available at https://github.com/CoderPowerBeyond/AutoHFormer.
@@ I Introduction
Time series forecasting [li2023revisiting, wang2024mamba, zhang2025fldmamba, huang2024long, wu2022timesnet] stands as a fundamental pillar of modern predictive analytics, enabling data-driven decision making across numerous mission-critical domains. As demonstrated in recent literature [yi2018integrated, zhangprompts], this task has become increasingly vital in our data-rich era. In financial markets [greenwood1997financial, zhang2024survey], accurate forecasting drives risk management strategies. Energy sector applications require precise demand predictions to optimize renewable integration. Transportation systems [yuan2021effective, zhang2025efficient, zhang2025survey] rely on temporal pattern analysis for traffic flow optimization. The importance of time-series analysis also extends to public safety through crime prediction [yi2018integrated] and model user behaviors with time evolving [zhang2025hmamba, zhang2025m2rec]. These diverse applications share a common need: the ability to extract meaningful patterns from temporal data to inform proactive decision-making.
@@
The evolution of time series forecasting has been significantly advanced by Transformer architectures [vaswani2017attention], yet persistent limitations hinder their practical deployment. As shown in Figure 1, conventional Transformers suffer from anti-causal attention flows, violating the fundamental $p(x_{t}|x_{<t})$ principle through bidirectional information propagation, which undermines autoregressive modeling. RNN-Transformer hybrids, while addressing causality, impose sequential computation bottlenecks ($\mathcal{O}(L)$ time steps, where $L$ is the sequence length), negating the parallel processing advantages of Transformers. Three fundamental shortcomings emerge from this architectural landscape: (1) Quadratic Complexity: Standard self-attention’s $\mathcal{O}(L^{2})$ memory and computation requirements make long-sequence processing infeasible; (2) Error Propagation: Autoregressive methods accumulate prediction errors exponentially across time steps; (3) Temporal Representation Rigidity: Fixed attention patterns fail to adapt to varying time scales, leading to suboptimal performance on multi-scale patterns.
Recent Advances and Limitations: Attention mechanisms have emerged as a powerful framework for temporal modeling, aiming to simultaneously optimize three key objectives: computational efficiency, temporal coherence, and multi-scale pattern capture. While recent innovations such as Informer [zhou2021informer], Pyraformer [liu2021pyraformer], and PatchTST [huang2024long] address individual aspects, they fail to achieve all three objectives simultaneously. Informer employs sparse attention for scalability but sacrifices multi-scale modeling; Pyraformer enforces hierarchical pyramidal structures that incur $O(L\log L)$ complexity, limiting adaptability to dynamic temporal scales; PatchTST efficiently handles long sequences but relies on anti-causal patching that violates strict causality. Similarly, GNN-based methods [chen2024graph, cirstea2021graph] model variable dependencies but struggle with non-stationary patterns and computational efficiency for long-range dependencies. These partial solutions leave the three core challenges—causality, efficiency, and multi-scale adaptability—unresolved within a unified framework.
Our Contribution: Unlike existing approaches that address these challenges in isolation, our proposed method, AutoHFormer, introduces a cohesive framework that integrates Dynamic Causal Windows for efficient computation and strict temporal causality. By leveraging learnable receptive fields and position-aware exponential decay, AutoHFormer adapts to varying time scales while maintaining computational efficiency with $O(LW)$ complexity. This mechanism ensures selective focus on relevant temporal patterns, enabling multi-scale adaptability from minute-level transients to daily periodicities. Furthermore, our hierarchical attention framework unifies top-down segment-level attention with bottom-up step-level refinement, balancing coarse-grained temporal trends with fine-grained details. Unlike prior methods, AutoHFormer simultaneously achieves sub-quadratic scaling, strict causality, and robust long-horizon forecasting, addressing the three competing objectives in a unified manner. As validated by our empirical results across diverse benchmarks, AutoHFormer delivers state-of-the-art performance with improved stability, reduced error accumulation, and no computational bottlenecks.
@@
  • A Unified and Efficient Hierarchical Autoregressive Framework for Time Series Forecasting. We propose AutoHFormer, a novel hierarchical autoregressive Transformer that integrates block-level (segment) processing with step-level refinement. This design balances coarse-grained temporal trends with fine-grained details, enabling efficient and accurate long-horizon forecasting.
  • Dynamic Causal Attention with Temporal Coherence. We introduce a memory-efficient attention mechanism based on Dynamic Causal Windows, which combines learnable receptive fields with position-aware exponential decay. This mechanism enforces strict causality while maintaining temporal coherence and achieving sub-quadratic complexity ($O(LW)$).
  • Adaptive Temporal Encoding for Multi-scale Patterns. To address varying temporal scales, we propose a relative positional encoding method with learned decay coefficients. This approach captures both short transients and long-term multi-scale patterns, enabling robust performance across diverse datasets and scenarios.
  • Comprehensive Empirical Validation. We conduct extensive experiments on benchmark datasets, demonstrating that AutoHFormer outperforms 11 state-of-the-art baselines, including transformer variants (iTransformer, Informer, Autoformer) and modern architectures (PatchTST, TimeMixer). AutoHFormer achieves particularly strong results in long-sequence forecasting (720+ prediction steps). All implementations, including hierarchical autoregressive mechanisms for our method and baselines, are available at https://github.com/CoderPowerBeyond/AutoHFormer.
@@ II-A Overview of AutoHFormer
[…]
@@ II-B2 Problem Definition
$\widehat{y}_{(h-1)H+t}=\mathcal{G}\big(\mathbf{X}_{1:L},\widehat{\mathbf{Y}}_{1:h-1},\widehat{y}_{(h-1)H+1},$., (3)
$\widehat{y}_{(h-1)H+t-1}\big)\quad\forall t\in\{1,.,H\}$
@@ II-C Hierarchical Autoregressive Generative Mechanism
    $\widehat{y}_{(h-1)H+t}=\mathcal{G}_{\phi}(\mathbf{C}_{h},\widehat{y}_{(h-1)H+1},.,\widehat{y}_{(h-1)H+t-1})$ (7)
    $\quad\forall t\in\{1,.,H\}$
@@
    $\mathbf{o}_{t}$ $=\text{WindowedAttention}(\mathbf{C}_{h}^{t},\mathbf{C}_{h}^{t},\mathbf{C}_{h}^{t};A)$ (8)
@@ II-D1 Dynamic Windowed Masked Attention (DWMA)
Time series forecasting demands efficient modeling of hierarchical temporal patterns, where both fine-grained fluctuations and coarse-grained trends jointly determine future states. Existing attention mechanisms [vaswani2017attention] fundamentally struggle with this multi-scale requirement due to either quadratic complexity ($\mathcal{O}(L^{2})$) or information loss in simplified approximations. Our DWMA breakthrough addresses these limitations via: i) Adaptively Constrained Attention Windows: The designation of adaptively constrained attention window is shown as $\mathcal{W}_{t}=\left[\max\left(1,t-\frac{W}{2}\right),t\right]$, which establishes dynamic receptive fields that maintain strict causality while enabling $\mathcal{O}(LW)$ computational complexity. The window size $W$ acts as a hyperparameter controlling the trade-off between context range and efficiency. ii) Learnable Exponential Decay: It introduces position-sensitive weighting via $\tau(t,t^{\prime})=\exp\left(-\gamma|t-t^{\prime}|\right)$, where the trainable parameter $\gamma$ automatically adapts to dataset characteristics. This creates continuous attention spectra that emphasize recent patterns while preserving access to critical long-range dependencies.
DWMA’s hybrid design achieves superior efficiency-accuracy trade-offs compared to vanilla attention [vaswani2017attention]. The decay mechanism automatically adjusts to diverse temporal patterns, which enables our model to have better performance in time series datasets.
@@ II-D3 Obtaining Embeddings Based on Enhanced Attention
In former steps, the attention weights are computed by incorporating the Precomputed Position Encodings (PPE). For each time step $t$ and its neighboring time steps $t^{\prime}$ within the sliding window of size $W$, the attention weights $A_{t,t^{\prime}}$ are calculated as above Equation 11. The attention weights $A_{t,t^{\prime}}$ are then used to compute the hidden states $\mathbf{H}_{\text{transformer}}$ for each time step $t$: $\mathbf{H}_{\text{transformer},t}=\sum_{t^{\prime}=1}^{L}A_{t,t^{\prime}}\cdot V_{t^{\prime}}$. Here $V_{t^{\prime}}\in\mathbb{R}^{d_{v}}$ is the value vector for time step $t^{\prime}$ and $d_{v}$ is the value dimension. The resulting hidden states $\mathbf{H}_{\text{transformer}}\in\mathbb{R}^{L\times d}$ capture the temporal dependencies in the input time series data, enriched by the relative position encodings.
@@ II-E Model Optimization
The training objective combines prediction accuracy with architectural constraints tailored for autoregressive time series modeling (Algorithm 1) via Equation 4. Adam updates with gradient clipping (max norm 1.0) and learning rate $\eta=10^{-4}$. The moving average update minimizes error accumulation through $\widehat{y}_{(h-1)H+t}\leftarrow W_{o}\mathbf{f}_{t}[-1]$.
@@ II-F Model Complexity
[…]
@@
[…]
@@ II-G Algorithm of AutoHFormer
[…]
@@
[…]
9:  $\mathbf{M}_{h}\leftarrow\text{SegmentMask}(\mathbf{C}_{h},H)$ $\triangleright$ Segment causality mask
@@
11:  $\widehat{\mathbf{Y}}_{h}^{init}\leftarrow\mathcal{F}_{\theta}(\mathbf{C}_{h})\in\mathbb{R}^{H\times d}$ $\triangleright$ Initial segment prediction
@@
15:   $\mathcal{W}_{t}\leftarrow\{t^{\prime}|\max(1,t-W/2)\leq t^{\prime}\leq t\}$ $\triangleright$ Local causal window
@@
17:   $\tau_{t}\leftarrow\exp(-\gamma\cdot|t-\mathcal{W}_{t}|)$ $\triangleright$ Adaptive decay
@@
22:   $\widehat{y}_{(h-1)H+t}\leftarrow W_{o}\mathbf{f}_{t}[-1]$ $\triangleright$ Final prediction
@@
24:  $\widehat{\mathbf{Y}}\leftarrow\text{Concat}(\widehat{\mathbf{Y}},\widehat{\mathbf{Y}}_{h})$ $\triangleright$ Aggregate segments